# Accelerating AllReduce with a Persistent Straggler

### v1

@@ Abstract
Distributed machine learning workloads use data and tensor parallelism for training and inference, both of which rely on the AllReduce collective to synchronize gradients or activations. However, bulk-synchronous AllReduce algorithms can be delayed by a persistent straggler that is slower to reach the synchronization barrier required to begin the collective. To address this challenge, we propose StragglAR: an AllReduce algorithm that accelerates distributed training and inference in the presence of persistent stragglers. StragglAR implements a ReduceScatter among the remaining GPUs during the straggler-induced delay, and then executes a novel collective algorithm to complete the AllReduce once the straggler reaches the synchronization barrier. StragglAR achieves a 2×2$\times$2 × theoretical speedup over popular bandwidth-efficient AllReduce algorithms (e.g., Ring) for large GPU clusters with persistent stragglers. On an 8-GPU server, our implementation of StragglAR yields a 22% speedup over state-of-the-art AllReduce algorithms.
@@ 1 Introduction
Distributed training and inference rely on collective communication primitives to transfer model gradients and activations between multiple GPUs. For example, in data parallelism, each GPU computes local gradients based on its shard of the training data and then averages its gradients with those computed by other GPUs to update the model. This gradient reduction across GPUs is achieved using the AllReduce collective communication primitive. AllReduce, like other communication primitives, is implemented in collective communication libraries (CCLs) like Nvidia’s NCCL.
Synchronous collective operations. Widely adopted CCLs implement bulk synchronous collective communication. In the bulk-synchronous model, every GPU in the communication group waits for all others before beginning the collective operation. This synchronization is achieved by invoking a synchronization barrier before the collective operation, which enables the use of highly optimized parallel algorithms to implement the collective communication operation.
Stragglers in synchronization delay distributed ML. Synchronization barriers are susceptible to straggler delays, caused when one or more GPUs in the communication group take significantly longer to initiate the AllReduce operation than the other GPUs. Figure 1 (left) shows how the straggler GPU (rank 0) initiates the collective operation late, causing all other GPUs to wait at the synchronization barrier before starting the collective. This is problematic in large-scale distributed ML, where the number of GPUs can be in the tens to hundreds and any one of them can delay the rest. In fact, straggler delays are known to reduce the throughput of large-scale ML jobs [1, 2].
Limitations of existing straggler mitigation techniques. Stragglers arise from a variety of runtime conditions (e.g., host-induced delays, network delays) or hardware-level inequalities between GPUs. To mitigate the impact of stragglers on collective communication, existing solutions either identify and remove straggler GPUs from ML clusters or approximate/drop the gradients delayed by stragglers. However, these solutions are not always effective. For example, removing straggler GPUs from the cluster leads to compute under-utilization. Estimating or dropping the straggler’s gradients can impact convergence in model training and does not generalize to tensor-parallel inference and training.
1.1 Our Proposal: Straggler-Aware AllReduce Algorithm with Provable Guarantees
We propose StragglAR, an algorithm that mitigates the impact of stragglers on AllReduce collective communication when a persistent straggler exists. A persistent straggler is a GPU that consistently takes longer to begin the AllReduce than other GPUs in the group. Persistent stragglers have been observed in our experiments (§2) and in large ML clusters [2]. Unlike prior straggler-mitigation methods [3, 4, 5], StragglAR does not modify data and is agnostic to the AllReduce use case, so it does not affect model convergence and applies to both data and tensor parallelism.
StragglAR leverages the delay caused by the straggler to perform useful communication among the other GPUs. Specifically, it implements a ReduceScatter among the other GPUs while the straggler is delayed. However, merely overlapping part of the communication with the straggler delay is insufficient. The newfound asymmetry between straggler and non-straggler GPUs makes it difficult to implement highly parallel communication that outperforms today’s heavily optimized algorithms. Thus, we design a collective algorithm that exploits the asymmetry to perform faster AllReduce at scale, achieving lower communication complexity by heavily parallelizing data transfers.
Assumptions. StragglAR improves the runtime of AllReduce under the following conditions:
  1. 1.
    The straggler is persistent, meaning that it is known which GPU rank consistently takes longer to finish computation and reach the synchronization barrier than the other GPUs.
  2. 2.
    The straggler is sufficiently delayed, typically on the order of a few milliseconds. The exact delay requirement depends on the AllReduce buffer size and is discussed in depth in §3 and §4.
  3. 3.
    The inter-GPU network enables any-to-any connectivity with homogeneous bandwidth between GPUs. This assumption is standard for contemporary multi-GPU servers and scale-up domains.
StragglAR outperforms the state-of-the-art. We verify that StragglAR can significantly reduce the overall communication time of AllReduce with experiments on multi-GPU servers and in simulation at larger scales. Our results on multi-GPU servers show that StragglAR can achieve a 22% speedup over popular AllReduce algorithms (e.g., Ring) in the presence of persistent stragglers. Our code is available at https://github.com/arjundevraj/stragglar.
2 Background & Motivation
In distributed training and inference, activations/gradients at each GPU are accumulated, reduced, and transferred over the network between GPUs using collective communication primitives, like AllReduce [6]. The parallelization strategy of an ML model determines which GPUs communicate and how often. Specifically, data parallelism and tensor model parallelism [7] use the AllReduce collective (Figure 2) to aggregate gradients and intermediate data, respectively, from multiple GPUs.
In ML software stacks, collective communication libraries (CCLs) implement algorithms for collective communication primitives (e.g., Nvidia CCL or NCCL). CCLs interface with ML application libraries (e.g., PyTorch [8], TensorFlow [9]) at one end and network transport protocols at the other end of the stack. Thus, the CCL decides how a group of GPUs exchanges data to implement a collective operation efficiently. For AllReduce, NCCL chooses between multiple algorithms based on the number of GPUs and the amount of data to be reduced (i.e., buffer size).
@@ 2 Background & Motivation
Most CCLs, including NCCL, adopt a bulk-synchronous model, where all GPUs in a communicator synchronize before the collective. This model is simple to reason about and enables the use of efficient parallel communication algorithms that rely on all ranks to implement the collective concurrently. However, it can significantly reduce performance when one or more GPUs in the communicator are slow to reach the synchronization point. The slowest GPU in the communicator is called the straggler, and the delays it causes are called straggler delays. Straggler delays can be due to hardware issues like thermal throttling and power supply, or runtime issues like network congestion, background noise, and skewed compute allocations across GPUs [1, 2, 10, 11]. In the synchronous model, tail effects mean that even with multiple slower GPUs, the slowest GPU (straggler) has an outsized impact on performance [3, 12, 13, 14, 15]. We explore related work about collective communication in distributed ML and straggler GPUs in §A.
While recent work [2] has shown persistent stragglers in ML jobs with thousands of GPUs, we show that even smaller multi-GPU jobs encounter persistent stragglers. We run Llama-3.2 fine-tuning jobs in three hardware environments: (1) 4 A100s with 40GB memory and NVLink 3.0 fully-connected mesh network in the Perlmutter supercomputer [16], (2) 4 A100s with 80GB memory in RunPod [17], and (3) 8 A100s with 80GB memory in RunPod. Both RunPod servers use NVLink 3.0 with an NVSwitched any-to-any network. For every environment, we repeat the same workload three times.
We encounter stragglers in all nine runs. We define the straggler delay as $T_{delay}=T_{rest}-T_{straggler}$, where $T_{straggler}$ is the AllReduce execution time of the straggler — the GPU that started the AllReduce last (rank 1 in Fig. 1) — and $T_{rest}$ is the AllReduce execution time of the second-slowest rank. Note that $T_{straggler}$ is the shortest AllReduce time among all GPUs because the straggler spends no time waiting for others before communicating, and our measure of $T_{delay}$ is conservative since it only encodes the idle time of the second-slowest GPU. We find that up to 50% of all iterations experience a straggler delay longer than 8 ms. The straggler GPU rank persists not only across iterations, but also across runs. Fig. 3 shows that one specific GPU is the slowest in a communication group in 98% of iterations, and is also the persistent straggler over multiple runs of the same workload. Fig. 3(a) shows these findings also generalize to multiple stragglers. Our experiments document (1) the severity of straggler GPUs that consistently delay AllReduce, and (2) that the same GPU rank is typically the straggler (i.e., persistent straggler).
(a) Perlmutter, 4 A100 SXM, Llama-3.2-1B, batch_size=32.
(b) RunPod, 4 A100 SXM, Llama-3.2-3B, batch_size=32.
(c) RunPod, 8 A100 SXM, Llama-3.2-3B, batch_size=64.
@@ 3 StragglAR: A Straggler-Aware AllReduce Algorithm
In this section we introduce StragglAR, a novel collective algorithm that accelerates AllReduce performance—with provably lower communication complexity—when there is a persistent straggler.
@@
$\alpha$-$\beta$ cost model. We reason about the performance of StragglAR using the $\alpha$-$\beta$ model of collective communication [18] widely used in collective algorithm design [19, 20, 21]. As per this model, sending a message of $s$ bytes takes $\alpha+s\beta$ time, where $\alpha$ is the fixed startup cost per message (hence independent of message size) and $\beta=\frac{1}{bandwidth}$ is the per-byte transmission cost in terms of time.
@@
Collective algorithms perform a series of data transfers among a set of GPUs to achieve the desired operation. These algorithms are typically specified as a schedule that describes the data transferred between a pair of GPUs over discrete time steps, where each time step is called a round. The $\alpha$ cost of an algorithm corresponds to the number of rounds since each round does at least one data transfer, and the $\beta$ cost is the total number of sequential bytes sent (number of rounds $\times$ bytes sent per round).
@@
The goal of a collective algorithm is to provide a data transfer schedule that minimizes the total $\alpha{-}\beta$ cost. Since $\alpha$ and $\beta$ are typically trade-offs, CCL implementations use $\alpha$-efficient algorithms for smaller buffers, when latency dominates, and $\beta$-efficient algorithms for larger buffers, when bandwidth becomes the bottleneck (since only $\beta$ cost scales with the buffer size). Different algorithms impact the coefficients on $\alpha$ and $\beta$ while the actual values of $\alpha$ and $\beta$ are treated as constants of the hardware. Sophisticated AllReduce algorithms bound the $\beta$ cost at ${\sim}2s\beta$ regardless of the number of GPUs through heavily parallelized communication.  StragglAR—like typical algorithms used for AllReduce in distributed ML (e.g., Ring)—focuses on optimizing $\beta$ cost because today’s large model sizes require communicating large buffers of several MBs to GBs.
@@
Key insights. When traditional AllReduce algorithms execute on multi-GPU setups, a straggler GPU can make all other GPUs wait until it is ready to begin the collective (GPU 0 in Fig. 1). In contrast, StragglAR uses this delay to perform useful work, by performing a ReduceScatter among non-straggler GPU ranks, until the straggler becomes available. By productively using the straggler’s delay, StragglAR reduces the data transfer cost once the straggler begins to communicate.
@@
Once the ReduceScatter is complete, StragglAR initiates a schedule of data transfers among all GPUs (both the straggler and non-stragglers) to complete the AllReduce (Fig. 2). However, even with the precondition, designing a faster algorithm than $\beta$-efficient baselines is challenging because the precondition introduces inherent asymmetry (straggler vs. other ranks) that makes communication more difficult to parallelize. Thus, StragglAR carefully designs communication to maximize parallelism in every round of the algorithm, achieving schedules that complete AllReduce faster than baselines. Further, computing efficient data transfer schedules is known to be combinatorially explosive [20], so StragglAR leverages symmetry-breaking to quickly find efficient schedules in polynomial time. For the desired multi-GPU setup, StragglAR is run once offline to generate AllReduce schedules, which are then implemented to run in real-time without modification. Next, we discuss the exact details of the runtime environment StragglAR relies on to improve AllReduce.
Identifying the persistent straggler. StragglAR must know the persistent straggler’s rank, which can be done with profiling tools such as nsys [22]. While we do not assume a specific lower bound on the straggler’s delay time, ideally it is long enough to fully overlap with the initial ReduceScatter. Even if the delay is shorter than this ideal, StragglAR often outperforms baselines, as shown in §4.
GPU cluster topology. Inter-GPU networks typically fall into two categories: scale-up (high-bandwidth connections within a node or rack) and scale-out (lower-bandwidth connections spanning multiple nodes or racks). Traditional AllReduce algorithms (e.g., Ring, recursive halving/doubling [19]) target the scale-up domain due to its homogeneous, any-to-any topology. In contrast, algorithms for scale-out environments rely on manual heuristics [20], specialized topologies (e.g., rail-optimized [23]), or hierarchical approaches layered on scale-up algorithms. As a result, StragglAR specifically addresses the scale-up domain, which can have 10-100s of GPUs [24].
@@
GPU-GPU connectivity. We assume each GPU has a single connection to the inter-GPU network, as is typical in scale-up domains [25]. Thus, at a given time, a GPU can only send data to one other GPU. While multidimensional topologies [26] enable using multiple links in parallel, this reduces the available per-link bandwidth in typical NVSwitched configurations. This bandwidth splitting increases the value of $\beta$ ($=\frac{1}{bandwidth}$), negating potential benefits from this form of parallelism.
@@
Input : $n=2^{k}$ GPUs, with rank $\sigma=n{-}1$ as the persistent straggler
@@
1 Initialization: rank $g<n{-}1$ holds partially reduced chunk $c_{g}$; straggler $\sigma$ holds none.
2 $A$: {}, dictionary of active chunks to set of GPUs that hold it

### v2

# Efficient AllReduce with Stragglers

@@ Abstract
Distributed machine learning workloads use data and tensor parallelism for training and inference, both of which rely on the AllReduce collective to synchronize gradients or activations. However, AllReduce algorithms are delayed by the slowest GPU to reach the synchronization barrier before the collective (i.e., the straggler). To address this challenge, we propose StragglAR: a parallel algorithm for AllReduce that accelerates distributed training and inference by exploiting natural variation in GPU execution times. StragglAR implements a ReduceScatter among the remaining GPUs during the straggler-induced delay, and then executes a novel collective algorithm to complete the AllReduce once the final GPU reaches the synchronization barrier. StragglAR achieves a $2\times$ theoretical speedup over popular bandwidth-efficient algorithms for large GPU clusters, surpassing the lower bound for bandwidth-optimal synchronous AllReduce by leveraging the asymmetry in when GPUs reach the synchronization barrier. On an 8-GPU server, StragglAR provides a 25% speedup over state-of-the-art AllReduce algorithms.
@@ 1 Introduction
Distributed training and inference rely on collective communication primitives to exchange model gradients and activations across multiple GPUs. In particular, the AllReduce primitive is used to average gradients across GPUs during data-parallel training and aggregate partial activations in tensor-parallel training and inference. AllReduce and other communication primitives are implemented in collective communication libraries (CCLs), like NVIDIA’s NCCL, which provide the core communication infrastructure for distributed ML.
To enable efficient communication, CCLs implement AllReduce using optimized parallel algorithms rooted in decades of research in high-performance computing [Dongarra et al., 2013]. These algorithms are designed to minimize communication time across GPUs — a key bottleneck in scaling modern ML workloads with massive data transfer sizes. To meet this demand, CCLs use bandwidth-optimal algorithms (e.g., Ring, Recursive Halving/Doubling), which heavily parallelize communication to fully leverage available inter-GPU network bandwidth.
Our key insight. Today’s collective algorithms are built on a fundamental assumption: all GPUs initiate the collective simultaneously. However, in distributed ML, communication depends on preceding computation; the last GPU to finish the preceding computation, i.e., the straggler, delays the entire AllReduce (Fig. 1, left). Our experiments show that this happens regularly even within a scale-up domain — the high-performance multi-GPU servers (e.g., NVIDIA DGX) and racks (e.g., NVIDIA GB200) that are the workhorses of distributed ML. In our experiments fine-tuning Llama-3.2 (1B, 3B) within DGX servers, we observe straggler delays of up to 30 milliseconds (Fig. 2(a)), prolonging the AllReduce execution time and causing other GPUs to stall. While stragglers in large datacenters are associated with network delays and hardware faults [Wu et al., 2024, Lin et al., 2025, Warraich et al., 2025, Jiang et al., 2024], we observe stragglers as an inherent challenge in distributed computation, where execution times naturally vary across GPUs. Existing mitigation strategies that approximate or drop the straggler’s data can impact model convergence and do not generalize to AllReduce in tensor-parallel training/inference. Thus, there remains a pressing need for bandwidth-efficient AllReduce that is robust to stragglers while still preserving correctness.
(a) Straggler delays in Llama-3.2 fine-tuning jobs.
(b) Range of simulated end-to-end speedups.
Figure 2: (a) CDFs of the straggler delay—the time between when the slowest and second-slowest ranks initiate the AllReduce—in Llama-3.2 (1B, 3B) fine-tuning jobs on the Perlmutter supercomputer and RunPod VMs (with 3 independent runs per job). (b) End-to-end speedups for ML workloads using StragglAR compared to bandwidth-optimal algorithms (e.g., Ring) for a 4 GB buffer (data-parallel buffer size of Llama-3.2-3B with local batch size of 4). Results are simulated with the $\alpha{-}\beta$ model using empirical $\alpha$ and $\beta$ values for the NVIDIA H100 DGX. Exposed communication is the percent of end-to-end time spent on AllReduce. As cluster size scales, StragglAR shows large gains (+30%) while worst-case performance is on par with the baseline (-3%).
Our proposal. Instead of treating stragglers as an anomaly, we propose designing algorithms that expect and exploit them. We design a novel AllReduce algorithm, StragglAR (Fig. 1, right), that provably transmits up to $2\times$ fewer bytes than the known bandwidth-optimal lower bound by exploiting natural variation in GPU execution times. StragglAR leverages the straggler delay to perform useful communication, eagerly executing a ReduceScatter among the other GPUs. However, the resulting asymmetry between the data buffers of the straggler and non-straggler GPUs after the ReduceScatter introduces new challenges in implementing highly parallel communication that can outperform known bandwidth-optimal algorithms. Thus, we design a new parallel communication algorithm that utilizes this asymmetry to perform faster AllReduce, achieving provably lower communication complexity in straggler settings while closely matching the complexity of bandwidth-optimal algorithms even without stragglers (Fig. 2(b)). To our knowledge, our work is the first to show that the decades-old lower bound for bandwidth-optimal bulk-synchronous AllReduce [Rabenseifner, 2004, Thakur et al., 2005, Patarasuk and Yuan, 2009, De Sensi et al., 2024] can be surpassed by leveraging natural variation in GPU compute times, while still maintaining the correctness guarantees of synchronous communication.
[…]
2 Background and Related Work
In distributed ML, GPUs aggregate gradients and activations using the AllReduce collective communication primitive (Figure 3), which transmits and reduces data across the inter-GPU network. Data parallelism uses AllReduce to average gradients, while tensor parallelism invokes it many times per model pass to exchange activations [Shoeybi et al., 2019a]. Collective communication libraries (CCLs) like NCCL implement these primitives, choosing bandwidth-optimal algorithms (e.g., Ring) for the large buffer sizes common in modern ML workloads [Shah et al., 2023]. CCLs adopt a bulk-synchronous model where all GPUs must synchronize before the collective, but performance degrades when a straggler—the slowest GPU—delays synchronization [Warraich et al., 2025, Dean and Barroso, 2013, Wang et al., 2024a, Gangidi et al., 2024, Li et al., 2014]. While stragglers result from inherent heterogeneity in GPU execution times, severe stragglers can stem from hardware issues (thermal throttling, power supply) or runtime factors (network congestion, compute skew) [Wu et al., 2024, Jiang et al., 2024, Grattafiori et al., 2024, Xiong et al., 2024]. Recent work has highlighted the acute impact of stragglers at datacenter scale [Jiang et al., 2024, Lin et al., 2025], and our experiments show intrinsic 30 ms delays even within multi-GPU servers (Fig. 2(a)).
@@ 2 Background and Related Work
Related Work. Prior straggler mitigation strategies either identify and remove stragglers [Jiang et al., 2024, Lin et al., 2025], wasting compute and only addressing severe cases, or approximate/drop the straggler’s data [Warraich et al., 2025, Harlap et al., 2016, Karakus et al., 2017, Recht et al., 2011], limiting applicability to data-parallel training and affecting convergence. Systems approaches [Wu et al., 2024, Zhao et al., 2024a] adapt workload placement or select among known algorithms, e.g., AdapCC [Zhao et al., 2024a] uses known Tree algorithms that sacrifice bandwidth for latency. In contrast, StragglAR is a novel bandwidth-efficient AllReduce algorithm that preserves exact reductions for both data and tensor parallelism and treats stragglers as intrinsic to distributed computation. While recent works synthesize new collective algorithms for heterogeneous networks [Shah et al., 2023, Zhao et al., 2024a, Wang et al., 2020, Won et al., 2024, Zhang et al., 2024], in homogeneous scale-up domains these converge to classical bandwidth-optimal algorithms like Ring [Shah et al., 2023, Wang et al., 2020], which NCCL implements [Hu et al., 2025]. We provide an in-depth discussion of related work in Appendix A.
@@ 3 StragglAR: A Straggler-Aware AllReduce Algorithm
We now present StragglAR, a novel algorithm to speed up AllReduce in the presence of stragglers.
@@
$\alpha$-$\beta$ cost model. AllReduce algorithms are typically analyzed using the $\alpha{-}\beta$ model of collective communication [Hockney, 1994, Thakur et al., 2005, Shah et al., 2023, Won et al., 2023, Wang et al., 2025]. As per this model, sending a message of $s$ bytes takes $\alpha+s\beta$ time, where $\alpha$ is the fixed startup cost per message (independent of message size) and $\beta=\frac{1}{bandwidth}$ is the per-byte transmission cost.
@@
Collective algorithms perform a series of data transfers among GPUs to achieve the desired operation. These algorithms specify a schedule that dictates which GPUs exchange data in each discrete time step, called a round. The $\alpha$ cost of an algorithm corresponds to the number of rounds, and the $\beta$ cost corresponds to the total number of serialized bytes sent (number of rounds $\times$ bytes sent per round).
@@
The goal of a collective algorithm is to provide a schedule that minimizes the total $\alpha{-}\beta$ cost. CCLs use $\alpha$-optimal algorithms for small buffers, when latency dominates, and $\beta$-optimal algorithms for larger buffers, when bandwidth becomes the bottleneck (as only $\beta$ cost scales with buffer size). Different algorithms affect the coefficients on $\alpha$ and $\beta$ while the $\alpha$ and $\beta$ constants come from the hardware. Bandwidth-optimal algorithms (e.g., Ring) bound $\beta$ cost at ${\sim}2s\beta$ regardless of the number of GPUs. StragglAR, like standard AllReduce algorithms used in distributed ML, focuses on optimizing $\beta$ cost because today’s large models incur large buffers of many MB to several GB.
@@
Key insights. In standard AllReduce algorithms, all GPUs must wait for the final GPU, the straggler, before starting the collective (GPU 0 in Fig. 1). In contrast, StragglAR uses this delay to perform a ReduceScatter (Fig. 3) among non-straggler GPUs. By productively using the straggler’s delay, StragglAR reduces the data transfer cost once the straggler is ready.
@@
Once the ReduceScatter is complete, StragglAR initiates a schedule of data transfers among all GPUs (straggler and non-stragglers) to complete the AllReduce. However, even with the precondition, designing an algorithm faster than $\beta$-optimal baselines is challenging because the precondition introduces inherent asymmetry (straggler vs. other ranks) that is difficult to parallelize. Thus, StragglAR designs communication to maximize parallelism in every round, achieving faster AllReduce schedules than baselines. While computing efficient schedules is known to be combinatorially hard [Shah et al., 2023], StragglAR uses symmetry-breaking to quickly find them in polynomial time. For the desired multi-GPU setup, StragglAR is run once offline to generate AllReduce schedules, which are then implemented to run in real-time without modification.
StragglAR is agnostic to the cause of the straggler or detection method (§4). Instead, it is a fundamental parallel algorithm that communicates $2\times$ fewer bytes than the known lower bound for AllReduce during exposed communication in settings where overlap is possible, while retaining strong worst-case performance. We now describe the runtime environment StragglAR relies on.
GPU cluster topology. Like classical AllReduce algorithms (e.g., Ring, Recursive Halving/Doubling [Thakur et al., 2005]), StragglAR targets the scale-up domain—nowadays extending to 10-100s of GPUs [NVIDIA, 2024]—due to its homogeneous, any-to-any topology.
@@
GPU-GPU connectivity. We assume each GPU has a single connection to the inter-GPU network, as is typical in modern switched scale-up domains [NVLink & NVSwitch, ]. Thus, a GPU can fully utilize bandwidth by sending data to one peer at a time (confirmed empirically in §4). While multidimensional topologies [Jouppi et al., 2023] enable using multiple links in parallel, this proportionally reduces the available per-link bandwidth in typical NVSwitched configurations (as we show in §4).
@@
Input : $n=2^{k}$ GPUs, with rank $\sigma=n{-}1$ as the straggler (without loss of generality)
@@
1 Initialization: rank g<n1g<n{-}1 holds partially reduced chunk $c_{g}$; straggler $\sigma$ holds none.
2 AA: {}, dictionary of active chunks to the set of GPUs that hold it